OpenStudy (anonymous):

Tom, Linda and Alex have $120 dollars. Alex has the third of what Tom has and Linda has twice as much as Alex. How much money, in dollars, does Linda have?

OpenStudy (anonymous):

T+L+A=120 A=1/3 T or T=3A L=2A then: 3A+2A+A=120 solve it for A (Alex) & you'll find L (Linda)

OpenStudy (anonymous):

A=20 L=2A=40

OpenStudy (anonymous):

ok thank

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

if you are good with answer - pick a "Good Answer". it will remove this problem from Unanswered.